# Gradient Boosting Reinforcement Learning

### v1

@@ 4 Gradient Boosting Reinforcement Learning
Figure 2: The GBRL framework. The actor’s policy and critic’s value function are parameterized by $\theta_{k}$. For example, $\theta_{k}=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy. $\theta_{k}$ is calculated by summing all the outputs of trees in the ensemble. Starting from $\theta_{0}$, at each training iteration, GBRL collects a rollout and computes the gradient $\nabla_{\theta_{0}}J$. This gradient is then used to fit the next tree added to the ensemble, which is updated to $\theta_{1}$. This process repeats with each iteration fitting a new tree, refining the parameterization, and expanding the ensemble towards $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$, an approximated scaled sum of gradients with respect to past parametrizations.
In this work, we extend the framework of GBT to support AC algorithms in the task of RL. The objective in RL is to optimize the return $J$, the cumulative reward an agent receives. Unlike in supervised learning, the target predictions are unknown a priori. RL agents learn through trial and error. Good actions are reinforced by taking a step in the direction of the gradient $\nabla_{\pi}J$. This formulation aligns perfectly with functional gradient ascent; thus, in GBRL, we optimize the objective directly over the decision tree function class. This is achieved by iteratively growing the ensemble of trees $\{h_{i}\}$. The ensemble outputs $\theta$, representing AC parameters such as the policy $\pi$ and the value function. For example, $\theta=[\mu(\operatorname{\mathbf{s}}),\sigma(\operatorname{\mathbf{s}}),V(\operatorname{\mathbf{s}})]$ for a Gaussian policy. At each iteration, a new tree $h_{k}$, constructed to minimize the distance to $\nabla_{\theta_{k-1}}J$, is added to the ensemble. Here, The resulting method is an application of GBT as a functional gradient optimizer $\theta_{k}\approx\theta_{0}+\epsilon\sum^{k-1}_{m=0}\nabla_{\theta_{m}}J$.
However, RL presents unique challenges for GBT. RL involves a nonstationary state distribution and inherent online learning, causing gradients to vary in magnitude and direction. Large gradients in unfavorable directions risk destabilizing training or leading to catastrophic forgetting. Moreover, feedback in RL is provided through interactions with the environment and is not available a priori. This contrasts with supervised learning settings, where gradients decrease with boosting iterations, and targets are predefined. As a result, many of the key features that traditional GBT libraries rely on are not suitable. For example, GOSS [20], categorical feature encoding [36], early-stopping signals, pruning methods [53], and strategies to tackle online learning [55].
To address these challenges, we employ appropriate tools from the NN and GBT literature, such as batch learning [13, 40] to update the ensemble. At each boosting iteration, we fit a decision tree on a random batch sampled with replacement from the experience buffer. This approach helps handle non-stationary distributions and improve stability by focusing on different parts of the state space, allowing beneficial gradient directions to accumulate and minimizing the impact of detrimental ones. Additionally, GBRL fits gradients directly to optimize objectives, whereas traditional GBT methods require targets and need workarounds to utilize gradients effectively.
A common theme in AC algorithms is to utilize a shared approximation for the actor and the critic. We adopt this approach in GBRL, constructing trees where each leaf provides two predictions. GBRL predicts both the policy (distribution over actions) and the value estimate. The internal structure of the tree is shared, providing a single feature representation for both objectives and significantly reducing memory and computational bottlenecks. Accordingly, in GBRL we apply differentiated learning rates to the policy and value outputs during prediction, effectively optimizing distinct objectives within this shared structure. We present the full algorithm in Algorithm 1 and diagram in Figure 2.
Algorithm 1 Gradient Boosting for Reinforcement Learning (GBRL)
1:Initialize: $\theta_{0}$, $\epsilon_{\text{actor}}$, $\epsilon_{\text{critic}}$, experience buffer $\mathcal{B}$, total training iterations $K$, number of updates $U$, batch size $N$, $k\leftarrow 1$
2:while $k<K$ do
3:     Collect trajectory $\tau^{(k)}=(\operatorname{\mathbf{s}}_{0},\operatorname{\mathbf{a}}_{0},\ldots,\operatorname{\mathbf{s}}_{T},\operatorname{\mathbf{a}}_{T})^{(k)}$ and rewards $(\operatorname{\mathbf{r}}_{0},\ldots,\operatorname{\mathbf{r}}_{T})^{(k)}$ using $\pi_{\theta_{k-1}}$
4:     Add trajectory $\tau^{(k)}$ and rewards to the experience buffer $\mathcal{B}$
5:     for each update $u=1$ to $U$ do
6:         Sample a batch from the experience buffer $\mathcal{B}$
7:         Compute gradients $g$ according to AC algorithm (e.g., PPO, A2C, AWR)
8:         Construct dataset $D=\{(\operatorname{\mathbf{s}}_{n},g_{n})\}_{n=0}^{N}$ and fit a decision tree $h_{k}$
9:         for each dimension $d=0$ to $D$ do
10:              if $0\leq d<D$ then
11:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{actor}}h_{k}(\operatorname{\mathbf{s}})$
12:              else
13:                  Update $\theta^{(d)}_{k}=\theta^{(d)}_{k-1}+\epsilon_{\text{critic}}h_{k}(\operatorname{\mathbf{s}})$
14:         $k\leftarrow k+1$
15:Output: AC parameters $\theta^{(d)}_{K}(\operatorname{\mathbf{s}})$ for $d=0,1,\ldots,D$
5 Experiments
Our experiments aim to answer the following questions:
  1. 1.
    GBT as RL Function Approximator: Can GBT-based AC algorithms effectively solve complex high-dimensional RL tasks?
  2. 2.
    Comparison to NNs: How does GBRL compare with NN-based training in various RL algorithms?
  3. 3.
    Benefits in Categorical Domains: Do the benefits of GBT in supervised learning transfer to the realm of RL?
  4. 4.
    Comparison to Traditional GBT libraries: Can we use traditional GBT libraries instead of GBRL for RL tasks?
  5. 5.
    Evaluating the shared AC architecture: How does sharing the tree structure between the actor and the critic impact performance?
We implemented GBT-based versions of A2C, PPO, and AWR within Stable Baselines3. We refer to our implementations as PPO GBRL, A2C GBRL, and AWR GBRL. We evaluated GBRL against the equivalent NN implementations. Where available, we utilize hyperparameters from RL Baselines3 Zoo [38]; otherwise, we optimize the hyperparameters for specific environments. The AWR NN implementation is based on the original paper [35].
We conducted experiments on a range of RL domains. We test classic control tasks, high-dimensional vectorized problems, and finally categorical tasks. We use 5 random seeds per experiment on a single NVIDIA V100-32GB GPU. We present the cummulative non-discounted reward, averaged across the last 100 episodes. We normalize the plots for simple visual comparison between GBRL and the corresponding NN implementations. The normalized score is computed as $\text{score}_{\text{norm}}=\frac{\text{score}_{\text{GBRL}}-\text{score}_{\text{NN}}}{\text{score}_{\text{max}\{\text{NN},\text{GBRL}\}}-\text{score}_{\text{min}\{\text{NN},\text{GBRL}\}}}$. We provide the full learning curves, implementation details, compute resources, un-normalized numerical results, and hyperparameters in the supplementary material.
Classic Enviroments.
We evaluate GBRL’s ability to solve classic RL tasks using Continuous-Control and Box2D environments, provided via Gym [50]. We trained agents for 1M steps (1.5M for LunarLander-v2) and provide the results in Figure 3. For exact values, refer to Table 2.
Considering the algorithmic objective, we observe that GBRL and NN present similar performance when optimized using PPO. In contrast, the other methods demonstrate inconclusive results. In certain environments, such as MountainCar, GBRL outperforms NN with all AC methods. On the other hand, in Pendulum NN is better.
Figure 3: Continuous-Control and Box2D environments. Normalized comparison between GBRL and NN. PPO, the best performing method, shows similar performance with GBRL and NN function classes.
High-Dimensional Vectorized Environments.
The decision-tree function class operates on individual features at each step. Consequently, this function class is not well-suited for handling pixel-based representations, which require more complex feature interactions. Therefore, we evaluted GBRL in the Football [22] and Atari RAM [4] domains. These offer high-dimensional vectorized representations. We trained agents in both environments for 10 million timesteps. The complete results are reported in Tables 3 and 4 and illustrated in Figure 4.
The results portray the following phenomenon. While both tasks may seem similar, there is a distinct difference. The features in the football domain are manually constructed and represent identifiable information, such as the location of the ball and the players. However, the Atari RAM domain provides a flattened view of the system RAM, which is unstructured.
At their core, binary decision trees are if-else clauses. This function class is naturally suited to work with structured data. These insights are emphasized in the football domain. Here, PPO GBRL greatly outperforms PPO NN across most environments and exhibits equivalent performance on the rest. In addition, as Atari RAM is unstructured, we observe that, as can be expected, in most cases GBRL underperforms NN, except for AWR. However, AWR NN underperformed considerably compared to the other NN implementations.
(a) Football Academy environments.
(b) Atari-ramNoFrameskip-v4 environments.
Figure 4: High-Dimensional Vectorized Environments. GBRL outperforms NN on the structured Football domain using PPO. NN outperforms on unstructured tasks, such as Atari RAM.
Categorical Environments.
The football experiment suggests that GBRL outperforms when assigned structured data. Here, we focus on categorical environments. This is a regime where GBT excels in supervised learning. In these experiments, we evaluated the MiniGrid domain [8]. It consists of 2D grid worlds with goal-oriented tasks that require object interaction. We trained in PutNear, FourRooms, and Fetch tasks for 10M timesteps, matching the reported PPO NN in RL Baselines3 Zoo. We trained the remaining environments for 1M timesteps. We give the results in Figure 5. For exact numbers, see Table 5.
Figure 5: MiniGrid environments. GBRL combined with the PPO backend outperforms NN across a range of categorical environments.
In MiniGrid, GBRL outperforms or is on-par with NN in most tasks. Specifically, PPO GBRL is significantly better than PPO NN. We observe the same trend when comparing between environments. These results emphasize that GBRL is a strong candidate for problems characterized by structured data, specifically when using PPO as the algorithmic backend.
GBRL vs Traditional GBT Libraries.
Here, we compare GBRL with Catboost and XGBoost. We focus on the PPO variant. When comparing to the standard libraries, we utilize their built-in options for incremental learning, vectorized leaves, and custom loss functions. As both CatBoost and XGBoost do not support differential learning rates, we used separate ensembles for the actor and the critic. For the comparison, we use the CartPole-v1 environment, training for 1M steps. The results are shown in Figure 6.
As seen, standard GBT libraries are unable to solve RL tasks in a realistic timeframe. GBRL, however, efficiently solves the task while also remaining competitive with NN across a range of environments.
@@ 5 Experiments
Figure 6: Comparing to standard GBT libraries. CatBoost and XGBoost are intractable in RL. Specifically, CatBoost lacks GPU support for custom losses, leading to low FPS and early termination.
Evaluating the shared AC architecture.
Finally, we evaluated the benefits of using a shared AC architecture by training PPO GBRL on three MiniGrid environments. We train agents with shared and non-shared architectures for 10M timesteps and compare the score, GPU memory usage, and FPS. We provide the aggregated results in Figure 7, and environment-specific breakdowns in the supplementary.
The benefit of the shared structure is clear both in terms of GPU memory consumption and computation speed. By sharing the tree structure, GBRL requires less than half the memory and almost triples the training FPS. This is achieved without any negative performance on the resulting policy, as seen in the reward plot.
@@
Figure 7: Shared Actor-Critic. Sharing the tree structure significantly increases training efficiency and memory, without impacting on the score.
Result summary.
The performance of GBRL varied across RL algorithms, but environments like MiniGrid highlight the potential advantages of using GBT in RL. The results suggest that GBT’s strengths in handling structured and categorical data from supervised learning can effectively transfer to the RL domain. Conversely, GBRL underperformed in Atari-RAM environments, indicating that certain environments, characterized by unstructured observations, are less suited for GBTs.
The results can be explained by the findings of Grinsztajn et al. [15], which suggest that NNs have an inductive bias toward overly smooth solutions and that MLP-like architectures are not robust to uninformative features. The optimal solutions for Atari-RAM might be smoother, which could explain the better performance of NNs. On the other hand, McElfresh et al. [31] argue that GBT outperforms NNs on ‘irregular’ datasets. Tree-based models excel in handling irregular patterns and categorical data, aligning with GBRL’s success in environments like MiniGrid.
Comparing different algorithmic backbones, we find PPO to be the strongest. PPO GBRL excelled in the MiniGrid and Football domains, and performed comparably with NN in classic control tasks. PPO GBRL’s success can be attributed to its alignment with GBRL’s incremental learning strategy. On the other hand, A2C’s single gradient update per rollout may limit its effectiveness and contribute to its underwhelming performance in many environments. Similarly, AWR’s design for multiple sample updates results in very large ensembles, creating a trade-off between large, slow, and memory-intensive ensembles, and lighter, less performant versions.

### v2

@@ 4 Gradient Boosting Reinforcement Learning
Deep RL with NNs faces significant challenges when dealing with structured observations containing both numerical and categorical features. Similarly to tabular datasets in supervised learning, these observations consist of heterogeneous features representing fundamentally different types of information, such as numerical measurements alongside categorical attributes describing system states or task conditions. These features create highly irregular patterns, as they may independently affect the environment based on their context, scale, or statistical properties. NNs struggle to model such patterns effectively (Beyazit et al., 2023; Jeffares et al., 2024), and their performance heavily depends on preprocessing methods (Hancock & Khoshgoftaar, 2020; Gorishniy et al., 2022). In contrast, GBT naturally handles irregular patterns and structured feature spaces efficiently. To leverage this strength, we present GBRL, an adaptation of GBT to RL tasks with such representations that frequently arise in real-world problems.
4.1 GBT as Functional Gradient Descent
GBT (Friedman, 2001) combines decision tree ensembles with functional gradient descent to learn complex nonlinear functions (Mason et al., 1999). Given a loss function $L$ and a dataset, $D=\{(\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{N}$, where $\mathbf{x}_{i}\in\mathbb{R}^{M}$ and $\mathbf{y}_{i}\in\mathbb{R}^{D}$, GBT aims to learn a function $F$, mapping inputs $\mathbf{x}$ to outputs $\mathbf{y}$. Through functional gradient descent, GBT minimizes the expected loss $\mathbb{E}_{\mathbf{x},\mathbf{y}}[L(\mathbf{y},F(\mathbf{x}))]$ with respect to the learned function $F$. Unlike parametric gradient descent, which updates parameters, functional gradient descent updates a learned function directly.
Starting with an initial function $F_{0}$. At each iteration $k$, GBT refines $F_{k}$, by taking a step in the negative direction of the gradient $g_{k}:=\nabla_{F_{k}}L(\mathbf{y},F_{k}(\mathbf{x}))$. The negative gradient is projected onto the space of possible trees $h\in\mathcal{H}$ by constructing a new binary decision tree that minimizes the following objective:
$h_{k}=\arg\min_{h\in\mathcal{H}}\|-g_{k}-h(\mathbf{x})\|_{2}^{2}\,.$ (1)
Through this process, GBT minimizes the expected loss, resulting in the additive model:
$F_{K}(\mathbf{x})=F_{0}+\sum_{k=1}^{K-1}\epsilon h_{k}(\mathbf{x})\,,$ (2)
where, $\epsilon$ is the learning rate controlling the step size in each iteration.
While standard GBT frameworks excel at handling structured data in supervised tasks, they require a tailored solution for RL’s iterative, dynamic, and multi-objective nature.
4.2 The GBRL Framework
Library Incremental Learning (sec) Standard Learning (sec)
CatBoost $36.17\pm 2.77$ $1.15\pm 0.03$
XGBoost $67.11\pm 0.70$ $2.26\pm 3.70$
GBRL $\mathbf{5.47}\pm\mathbf{1.19}$ N/A
Table 1: Incremental Learning Speed in GBT Libraries. Training times (seconds) for 1000 boosting iterations on random batches of 128 samples with 20 input features and 3 targets. While CatBoost and XGBoost train efficiently in standard mode using the same batch for all iterations, their training time increases significantly in incremental learning, where each batch receives only one boosting iteration. GBRL, designed explicitly for incremental updates, remains much faster.
In the GBRL framework (Figure 1) we adapt GBTs to handle the challenges of RL. To do so we reinterpret $F_{K}(\operatorname{s})$ as the parameterization of the policy and value function, where the observed state $\operatorname{s}\in\mathcal{S}$ is the input to the ensemble.
Specifically, at each iteration $k$, we compute the policy gradient (Sutton et al., 1999). As opposed to the original formulation, where the gradient is calculated with respect to parameter space, here we calculate it in function space. Consequently, we express the policy gradient as:
$\nabla_{F_{k}}J(\pi_{F_{k}})=\mathbb{E}_{\pi_{F_{k}}}[\nabla_{F_{k}}\log\pi_{F_{k}}(\operatorname{a}|\operatorname{s})A(\operatorname{s},\operatorname{a})],$ (3)
where $F_{k}$ is the current ensemble-based parametrization.
Then, the policy gradient is projected onto a new tree and added to the ensemble:
$h_{k}=\arg\min_{h\in\mathcal{H}}\|\nabla_{F_{k}}J(\pi_{F_{k}})-h(\operatorname{s})\|_{2}^{2}\,.$ (4)
The result is an application of GBT as a functional gradient optimizer, updating both actor and critic incrementally:
$F_{K}(\operatorname{s})=F_{0}+\sum_{k=1}^{K-1}\epsilon h_{k}(\operatorname{s})\approx F_{0}+\sum^{K-1}_{k=1}\epsilon\nabla_{F_{k}}J(\pi_{F_{k}}).$ (5)
Gradient-based Approach.
As seen in Equation 5, GBRL is an online learning gradient-based optimizer. In each step, a new tree is constructed to minimize the sampled loss. This learning scheme is identical to the common policy gradient methods (Mnih et al., 2016). By leveraging optimization frameworks for gradient computation, such as PyTorch (Paszke et al., 2019), GBRL can be integrated with most Actor-Critic algorithms and implemented within existing RL libraries. In contrast, traditional GBT frameworks are designed for offline, static datasets and require significant customizations and workarounds to function in RL.
@@ 4.2 The GBRL Framework
Figure 2: GBT library comparison, Cartpole. CatBoost and XGBoost are intractable in RL. CatBoost’s lack of GPU support for custom losses leads to low FPS and early termination.
@@
Figure 3: Shared Actor-Critic, MiniGrid. The shared tree structure significantly increases efficiency, without impacting the score. Aggregated results over three tasks are shown here, full per-task curves are available in Appendix (Figure 20)
Shared Actor-Critic Architecture.
Many RL algorithms operate with a shared Actor-Critic structure. This structure helps avoid overfitting and improves overall convergence speeds (Henderson et al., 2018; Andrychowicz et al., 2020). We adopt this approach in GBRL such that each leaf produces two predictions. GBRL predicts both the policy (distribution over actions) and the value estimate. The internal structure of the tree is shared, providing a single feature representation for both objectives and significantly reduces memory and computational bottlenecks. Similar to the standard practice, we support and use separate learning rates for the policy and value. This approach enables solving two distinct objectives within a single shared structure.
4.3 Comparison to Traditional GBT Libraries
Before diving into the comparisons with NNs, Section 5, we first evaluated our design choices against traditional GBT libraries. While these libraries can be adapted for RL using custom loss functions and incremental learning, experiments with CatBoost and XGBoost reveal scalability issues inherent to their supervised learning design. Table 1 shows these frameworks are significantly slower than GBRL, when used for incremental learning. Incremental learning is a fundamental property of RL, as the policy is continually changing. Moreover, when training on CartPole (Figure 2), larger ensemble sizes lead to significant drops in training throughput. In contrast, our shared Actor-Critic architecture halves memory usage and doubles throughput without compromising policy quality (Figure 3). Overall, GBRL offers a purpose-built solution that maintains GBT’s advantages while efficiently handling RL-specific needs.
5 Experiments and Results
Our experiments aim to answer two core questions:
1. GBT as an RL function approximator: Can GBT-based AC algorithms effectively solve complex high-dimensional RL tasks? And how do they compare with NNs?
2. Advantages of GBT: Building on the GBT’s success in irregular and tabular data, does its inductive bias offer similar robustness benefits in RL – specifically for out-of-distribution states, noisy inputs, and spurious correlations?
For our experiments, we implemented a GBT-based version of PPO within Stable Baselines3 (Raffin et al., 2021). Where available, we use standard hyperparameters, environment-specific, and normalization wrappers according to RL Baselines3 Zoo (Raffin, 2020); otherwise, we optimize the hyperparameters for specific environments. For each experiment, we report the aggregated performance across five random seeds. Our training setup consists of a single NVIDIA V100 GPU. We refer the reader to the supplementary material for additional technical details, such as hyperparameters, implementation, and environment details (Appendix C), full results (Appendix D) and full learning curves (Appendix E) for all evaluated algorithms.
5.1 Standard Environments
Experimental Setup.
We evaluated GBRL against NNs across three categories of RL benchmarks. First, we tested classic RL tasks using Classic-Control and Box2D environments from Gymnasium (Towers et al., 2024). These provided a baseline for comparison on standard benchmarks. We then evaluated both methods on more complex environments with high-dimensional vectorized representations: the Football domain (Kurach et al., 2020), where we employed the built-in ’Checkpoints’ shaped reward, and the Atari RAM domain (Bellemare et al., 2013). Finally, we assessed performance on categorical environments, specifically targeting the MiniGrid domain (Chevalier-Boisvert et al., 2023) – a setting where GBTs have traditionally excelled in supervised learning. MiniGrid provides an ideal test environment through its 2D grid-world structure featuring goal-oriented tasks with discrete object interactions.
Results.
We report the cumulative non-discounted reward, averaged across the last 100 episodes in Figure 4. For a simple visual comparison between GBRL and NNs, we report the normalized score: $\operatorname{r}_{\text{norm}}=\frac{\operatorname{r}_{\text{GBRL}}-\operatorname{r}_{\text{NN}}}{\operatorname{r}_{\text{max}\{\text{NN},\text{GBRL}\}}-\operatorname{r}_{\text{min}\{\text{NN},\text{GBRL}\}}}$.
Figure 4: GBRL vs NN in standard environments (PPO). Aggregated mean and standard deviation of the normalized average reward for the final 100 episodes.
Decision trees, at their core, are if-else structures that excel in handling numerical and categorical data. In MiniGrid, which provides such environments, GBRL significantly outperformed NN. For example, in tasks like PutNear, FourRooms, and Fetch, GBRL consistently achieved higher rewards, highlighting its ability to exploit the structured nature of the environment effectively. This performance demonstrates that GBRL is particularly well-suited for problems where data can be neatly partitioned, aligning with GBT’s inherent design strengths. This suggests that structured environments represent an ideal use case for GBRL.
An additional domain characterized by structured representations is the Football domain, where we used a shaped reward. Here, features such as player positions and ball locations were designed to represent identifiable, interpretable information. In this domain, GBRL outperformed NN across most scenarios and exhibited equivalent performance in the rest. The structured nature of the Football domain aligns well with the strengths of GBT, effectively partitioning the feature space. Interestingly, while both Football and MiniGrid are structured, the Football domain is not sparse. Hence, GBRL’s strong performance in Football underscores its robustness and capacity to generalize well in structured, high-dimensional environments.
Unlike Football, the Atari RAM domain offers unstructured, high-dimensional representations by exposing raw system memory states. These flattened views lack explicit semantic structure, making them challenging for GBRL, which relies on single-feature splits. As a result, GBRL underperformed compared to NN in most Atari RAM tasks. This result highlights an important limitation of decision-tree-based approaches. They struggle to model implicit feature interactions, an inherent characteristic of unstructured data.
Finally, classic control tasks are simpler environments that rely on low-dimensional state spaces. In these tasks, GBRL demonstrated performance comparable to that of the NN counterpart. This suggests that decision tree-based models can match NNs in solving tasks where feature interactions are limited and the optimization landscape is smooth.
Overall, these comparisons indicate that GBT can serve as a strong function approximator in RL, matching or surpassing NN’s performance in certain environments.
5.2 Challenge Environments
Having established GBRL’s general capabilities, we now examine scenarios designed to test inherent properties of decision trees. These experiments focus on challenging conditions—sparse rewards, misleading correlations, and input perturbations—where NNs typically struggle.
5.2.1 Signal Dilution
Signal dilution occurs when rare events, such as sparse rewards, are statistically overwhelmed by more common non-event data during training (Shyalika et al., 2024; Zhao et al., 2018; He & Cheng, 2021; Tessler et al., 2017). In NNs with fixed architectures, shared weights inherently prioritize patterns from dominant classes or frequent signals, causing rare events to be ”averaged out” or under-represented in learned representations. This phenomenon is particularly problematic in RL and classification tasks with imbalanced datasets, where the model fails to retain or react to critical but infrequent signals.
Experimental Setup.
We compare GBRL with NNs in a linear equation variable isolation environment with sparse rewards, inspired by (Poesia et al., 2021). The agent observes the coefficients $(a,b,c)$ of a linear equation $ax+b=c$, and needs to isolate $x$, receiving a large reward only upon completion. Actions are multi-discrete, involving the selection of an operation, a digit $\in[0,9]$, and its sign. The optimal policy solves the simple task in 2 steps. We also evaluate two more complex variations with an additional variable $y$, where the optimal policy requires 3–4 steps. For these, we introduce intermediate goals with small rewards to guide the agent.
Results.
While simple for humans, these tasks remain challenging for NNs due to signal dilution (Figure 5). Although they can make some progress in the simple task ($ax+b=c$), they fail when the equation complexity increases, even when given intermediate rewards. In contrast, GBRL converges to and retains a stable (and optimal) policy throughout the process. GBTs naturally segment the decision space into explicit rule-based splits, preventing signal dilution while correctly preserving the rare but crucial reward signals. Consequently, we see that GBRL can preserve and exploit rare but critical reward signals—evidence that GBT-based methods are inherently less prone to signal dilution.
Figure 5: Signal dilution, variable isolation task. Mean and standard deviation of the average episodic reward during training. GBRL was trained for 15M training steps and NN for 30M. Episodes are terminated after 50 steps if not solved.
5.2.2 Spurious Correlations
Experimental Setup.
The agent is placed in a grid environment with three balls (red, green, and blue), and a red box. The ball locations are randomized. The agent observes an objective specifying which ball to pick up. We experiment with three box-placement configurations: adjacent to the target ball, adjacent to one of the non-target balls, and without a box in the room. This task tests whether the policy overfits spuriously correlated features. We then analyze the results using SHAP (Lundberg & Lee, 2017), a method based on Shapley values (ARROW et al., 1953) that measures the contribution of each feature to the final prediction of a model.
Results.
GBRL consistently outperforms NNs in all scenarios (Figure 6), demonstrating its ability to ignore irrelevant distractions, such as the box, and reliably reach the target regardless of its placement.
To understand GBRL’s performance, we analyzed the SHAP values for the agent’s surroundings and the mission feature when predicting the logit for picking up the target object (Figure 7). GBRL assigns the strongest positive SHAP value to the target object, followed by the mission string, aligning perfectly with the goal. Non-target objects receive the most negative SHAP values, suppressing distractions and reinforcing goal-directed behavior. In contrast, NNs assign negative SHAP values to the mission feature and rely heavily on the grid surrounding the target object. This indicates a reliance on spurious correlations and a lack of goal-oriented strategies.
These findings underscore GBRL’s ability to focus on task-relevant features and ignore deceptive cues, reinforcing our hypothesis that GBT’s inductive bias provides robustness against spurious correlations. Next, we consider how this robustness extends to other forms of state perturbations.
Figure 6: Spurious correlations, MiniGrid. Mean and standard deviation of the episodic reward. Plots outlined in green are in-distribution, while plots outlined in red are out-of-distribution.
(a) GBRL SHAP.
(b) NN SHAP.
Figure 7: Spurious correlations revealed by feature importance in MiniGrid. The visualization shows normalized SHAP values for the concluding action in a successful evaluation run, revealing the magnitude and direction of each feature’s impact on model predictions. In this scenario, a red box was positioned next to the target (red ball), and both models converged to a successful policy. GBRL prioritizes the mission, ignores the red box, and contrasts the blue ball. In contrast, the NN focuses on the cells surrounding the target, while ignoring the blue ball and contrasting the box and the mission.
5.2.3 Robustness to State Perturbations
When encountering an out-of-distribution (OOD) state, a GBT model will always output a result within the same range seen in training. In contrast, NNs may extrapolate and exhibit unexpected behavior (Xu et al., 2021; Meinke & Hein, 2020; Ulmer & Cinà, 2021). We consider three tasks. First, robustness to irrelevant information. Second, robustness to random noise. And finally, robustness to missing features. These experiments aim to test the inherent robustness properties of GBTs when encountering different types of noise during training.
Irrelevant Information.
We consider a similar fetch experiment as before (Section 5.2.2). Now, a new object is placed in the room at a random location. This object is not part of the mission and only interferes with navigation. The interfering object differs from the task objects in shape, color, or both.
Figure 8 illustrates how introducing a distractor impacts policy performance. When the environment is free of distracting objects, NNs are able to solve the task, converging to a stable and ideal solution. Surprisingly, although the purple-box object shares no features with the task objects (blue, green, and red balls), NNs exhibit a drop in performance but are able to solve the task in a subset of seeds. This performance degradation increases when the confusing object shares more features with the goal objects, such as purple-ball and red-box. In contrast, GBRL successfully and efficiently solves the task, regardless of the additional irrelevant object placed in the scene.
Figure 8: Robustness to irrelevant objects, MiniGrid. Mean and standard deviation of the average episodic reward during training. GBRL was trained for 5M training steps, and NN for 20M.
Random Noise.
In the classic-control environments (Section 5.1), we train agents in a noiseless setting. Then, to evaluate the robustness of the resulting policies, we add random zero-centered Gaussian noise to the input states as defined by: $\operatorname{s}_{t}=\operatorname{s}_{t}+\epsilon|\operatorname{s}_{t}|$.
The results show that these characteristics translate to practical outcomes (Figure 9). In classic-control environments, a trained GBT-based policy is much more robust to small variations in the state space compared to NNs. We attribute this performance to how GBTs partition the input space. This partitioning occurs on hard thresholds that are split according to single features. Although small perturbations may be sufficient to completely change decision outcomes in NNs (Lütjens et al., 2020; Carlini & Wagner, 2017), GBRL exhibits robustness to random additive noise.
Figure 9: Robustness to input perturbations, classic-control. Models were trained without perturbations and then evaluated on noisy observations. At each timestep, the observation was perturbed following $\operatorname{s}_{t}=\operatorname{s}_{t}+\epsilon|\operatorname{s}_{t}|$, where $\epsilon\sim\mathcal{N}(0,\text{var})$. We report the mean and standard deviation of episodic reward across 100 evaluation episodes for each $\text{Variance}\in[0,0.2]$.
Missing Features.
We modify scenarios in the football domain (Section 5.1) to support the random drop of a player from the opposing team. Policies are trained on both the original task and a modified variant. We then evaluate model robustness by testing how a policy trained against $N$ random players performs when faced with $N+1$, or how a policy trained against a full team behaves when a random opponent is removed.
The results show that GBRL outperforms NNs in OOD tasks (Figure 10), particularly in the 11 vs 11 academy scenario. Notably, GBRL maintains performance when evaluated against 10 opponents and can solve 11 vs. 11 even when trained on 10. In contrast, NNs struggle in all OOD cases. This suggests GBRL better isolates features that are essential to solving the task.
Figure 10: Out-of-distribution evaluation, football. Mean and standard deviation of episodic reward across 100 evaluation episodes. Plots outlined in green are in-distribution, while plots outlined in red are out-of-distribution.
6 Ablation Study
In this section, we examine the sensitivity of GBRL to key hyperparameters. We systematically changed one hyperparameter at a time while keeping the others fixed. For each experiment, we trained an agent in the CartPole-v1 environment for 1M timesteps using GBT-based PPO.
Learning Rate Variation Across Policy and Value Function Components.
We investigated the impact of the learning rate for both the policy and value function components. In these experiments, we held one learning rate constant while varying the other. Consistent with observations in NN training, we found that excessively high learning rates destabilized the learning process, whereas overly low rates led to significantly slower convergence (as illustrated in Figure 11).
Figure 11: Ablation: Learning rate variation across policy (left) and value function (right) components. Mean and standard deviation of the average episodic reward during training. Each subplot shows the effect of varying the learning rate for one component (policy or value function) while keeping the other fixed.
Limitations of Tree Depth.
We varied the maximum depth of the decision trees built at each training iteration and evaluated their impact on learning dynamics. Figure 12 demonstrates that deeper trees yield more accurate gradient approximations, leading to faster convergence, but also incur a higher computational cost per training step. In the paper, we selected a maximum tree depth of 4, balancing convergence speed and wall-clock time.
Figure 12: Ablation: Effect of tree depth on convergence. We varied the maximum depth of the decision trees used in the gradient-boosted models and evaluated their impact on learning dynamics. The plot shows the mean and standard deviation of the average episodic reward over time and the total runtime.
Batch Size Impact on Convergence.
We analyzed the impact of the batch size on performance. In Figure 13, we observe that batch size significantly impacts convergence. Specifically, smaller batches result in GBRL building more trees per rollout, improving adaptability. However, smaller batches also lead to noisier gradient estimates as a result of limited samples per constructed tree. Conversely, larger batches stabilize training by reducing variance through averaging within leaves, as more samples are utilized to construct each tree but build fewer trees per rollout. Hence, both excessively small and large batch sizes negatively impact performance.
Figure 13: Ablation: Batch size impact on convergence. Mean and standard deviation of the average episodic reward during training. We fix the rollout length to 2048 and vary the batch size. Smaller batches improve early learning but suffer from instability, while larger batches stabilize training at the cost of slower convergence.
Overall, these ablations highlight that GBRL’s performance is sensitive to training frequency, tree depth, and batch size—each of which governs the trade-off between convergence speed and stability. While this section focused on general hyperparameters, we further investigate how these trade-offs manifest across different RL algorithms in Appendix B.
7 Limitations and Future Directions
GBRL exhibits competitive performance and surprising robustness capabilities. However, despite these benefits, we highlight two important limitations. (i) Continuous generation of trees. As the policy improves through numerous updates, the size of the ensemble increases. This unbounded growth has implications for memory usage, computational efficiency, and the feasibility of online real-time adaptation. Future work may leverage tree redundancy to produce a more efficient and compact policy. For example, strategies for tree pruning, ensemble compression, or dynamically managing ensemble size. (ii) Off-policy continuous control. Methods like DDPG (Lillicrap et al., 2019) and SAC (Haarnoja et al., 2018) update the policy by differentiating the Q-estimator. As GBTs are not differentiable, new solutions are needed to incorporate them into these algorithms.